### v2

Abstract.
Security engineering, from security requirements engineering to the implementation of cryptographic protocols, is often supported by domain-specific languages (DSLs). Unfortunately, a lack of knowledge about these DSLs, such as which security aspects are addressed and when, hinders their effective use and further research. This systematic literature review examines 120 security-oriented DSLs based on six research questions concerning security aspects and goals, language-specific characteristics, integration into the software development lifecycle (SDLC), and effectiveness of the DSLs. We observe a high degree of fragmentation, which leads to opportunities for integration. We also need to improve the usability and evaluation of security DSLs.
domain specific languages, security, survey
ccs: Security and privacyccs: Software and its engineering Domain specific languages
1. Introduction
Security is a fundamental aspect to be considered in the engineering of software-intensive systems, as security breaches can have serious consequences. Consider an Electronic Health Records (EHR) system such as iTrust (Heckman et al., 2018; Meneely et al., 2021): It is important that sensitive patient information is not exposed to unauthorized entities, but it would be even worse if an attacker were able to manipulate medical records, potentially leading to incorrect and harmful treatments. Although there is a relatively high level of awareness among developers today about the security of the software-intensive systems they develop (Xie et al., 2011), successful attacks on such systems are still frequently reported, such as hospitals being forced to shut down due to an attack on their EHR systems (Singh and Sittig, 2016).
One of the main challenges in developing secure software-intensive systems is that security engineering spans the entire Software Development Life Cycle (SDLC) (McGraw, 2004). Security has to be considered with entirely different levels of abstraction throughout the entire SDLC and security compliance among all developed artifacts must be ensured (Peldszus, 2022). However, already properly planning or realizing security on a single artifact is non-trivial. Relevant security requirements are often implicit (Bartsch, 2011), which hardens to address these systematically. Furthermore, common development artifacts, such as source code that is usually written in general purpose programming languages, do not support easy reasoning about security but security experts should be given system representations that are tailored towards their tasks (Peldszus, 2022).
Contrary to general purpose programming languages, Domain-Specific Languages (DSLs) (Wasowski and Berger, 2023) are purpose-built to cater to the needs of a particular problem domain. While DSLs can encompass programming languages, they can also target other development phases, such as design-time system descriptions. Leveraging the inherent specificity of the target domain, DSLs empower developers to address problems and engineer solutions at the most appropriate level of abstraction. By providing a more intuitive representation of systems, DSLs simplify the design process. Moreover, DSLs often offer the capability to expose parameters that are challenging to express with general-purpose languages. Furthermore, DSLs are crucial for leveraging powerful supporting tools, enabling optimal design exploration, automated testing, and formal verification of specific properties.
Due to their power, domain specific languages have been proposed for facilitating a wide range of security engineering tasks. In general, information security refers to the practice of protecting information from unauthorized access, disclosure, modification, or destruction. It involves a set of measures and techniques designed to ensure the three core principles: confidentiality, integrity, and availability of information (Cawthra et al., 2020). Multiple goals can be derived from these core principles or are strongly related, including authentication, privacy and accountability. To counter the many possible attack vectors, which threaten security goals, a wide range of defensive techniques and building blocks exist at various system levels. For example, access policies form a fundamental component, establishing guidelines and controls to determine who is granted access to specific resources within an organization’s network. At a lower level, cryptography provides an existential toolkit to enforce security policies by preventing malicious parties from reading and manipulating sensitive data.
DSLs play a critical role in addressing the unique requirements of specific domains—including security in particular. The scientific community has contributed a variety of security DSLs, targeting different phases of the SDLC and different aspects of security. This broad security spectrum is reflected in the highly fragmented research on security DSLs, which limits the widespread adoption of security DSLs and hinders their development. This survey aims to provide a comprehensive overview of DSLs in the scientific literature, focusing on those with security-specific language features. Our goal is to identify connections and gaps, and to suggest possible combinations and integrations. In particular, there is currently a lack of work that provides a holistic analysis of DSLs across different security domains. We make a first attempt to organize previous research in the area of DSLs for security. The following research questions guide our investigation:
  1. RQ1:
    What Security DSLs have been presented in the scientific literature?
    While the scientific community has contributed numerous security DSLs, there is currently no systematic overview of them. We will explore the existing literature and focus on DSLs that address security concerns in software development to provide such a missing catalog of security DSLs.
  2. RQ2:
    What security aspects are targeted by the security DSLs?
    As discussed, security DSLs have been designed to address different aspects of security in terms of different security goals, attacker models, and defense mechanisms. In doing so, they can be problem-oriented, focusing on general security planning and the identification of potential threats, or solution-oriented, facilitating the realization of concrete defense mechanisms. To this end, we aim to capture these aspects in order to contextualize the security DSLs.
  3. RQ3:
    What phases of the development process are supported by the DSLs?
    In the context of developing secure software-intensive systems, security DSLs need to be integrated into the development process. Thereby, DSLs can be relevant for a single development phase, but also for several phases. Accordingly, we aim to identify which SDLC phases are relevant for each security DSL and which security aspects the DSLs focus on in which phases of the software development lifecycle. In particular, for DSLs that aim to support multiple phases, we are interested in which phase they are instantiated and in which phases they are used.
  4. RQ4:
    What types of DSLs exist, and how are they used?
    When designing a DSL, language engineers can choose between different types of DSLs with different characteristics. To provide an overview of the existing DSLs that can serve as a reference, we will also examine their general characteristics, such as whether they are stand-alone (external) DSLs or embedded in other established languages (internal DSL), such as a general-purpose programming language, and whether they are textual or visual DSLs. In addition, we will examine the tools associated with these DSLs to understand how they are used in practice.
  5. RQ5:
    What are the semantics of the DSLs?
    Based on the targeted security goals and the SDLC phases in which the DSLs are addressing these goals, we investigate how this is semantically realized. Specifically, we are interested in what mechanism brings the DSLs instances into effect. To this end, we aim at capturing how are the instances of the DSLs executed to achieve the corresponding security goals or what kind of transformation is provided by the backend to propagate and utilize the security properties specified using the DSLs. We will also explore how the DSLs are integrated into the development process.
  6. RQ6:
    How effective are the DSLs?
    Lastly, we are interested in how effective the DSLs are and how their effectiveness has been proofed or validated. Primary, this questions aims at (formal) security guarantees including properties like soundness, precision, and completeness. However, we also are interested in empirical data on the security DSLs and whether the DSLs have been evaluated on real-world user studies.
With this paper we hope to support multiple groups of readers. First, we want to give developers of secure systems an overview of DSLs that can support their work at different stages of the software development life cycle. Our work is also aimed at security researchers who want to leverage DSLs for their specific security domain and are looking for examples, best practices and universal solutions from existing security DSLs. Lastly we target security researchers working on language-based security, possibly across multiple security domains and attacker models, interested in integration, interaction and combination of DSLs.
2. Background
We first give a general introduction to domain-specific languages, and then discuss information security in the context of domains that need to be considered when developing secure systems.
2.1. Domain-Specific Languages
A DSL is a programming or specification language that is dedicated to a particular problem domain, problem representation, or solution technique (Mernik et al., 2005; Fowler, 2011; Wasowski and Berger, 2023). DSLs are designed to simplify the development process within a specific context or domain by providing abstractions and syntax that directly represent concepts and operations relevant to that domain. Usually, they hide unnecessary details and complexity present in general-purpose languages, allowing developers to focus on solving problems within their domain of expertise, such as secure software design. Because DSLs are tailored to the needs of a specific domain, they are more expressive, concise, and easier to use within that domain than general-purpose programming languages. To this end, DSLs can greatly improve productivity and code maintainability within specific domains by allowing developers to express solutions using terminology and abstractions familiar to domain experts. However, designing and implementing DSLs requires careful consideration of the domain requirements, the target audience, and the tradeoffs between expressiveness and ease of use.
2.2. Information Security
Information security aims to control the flow of sensitive information at various system levels (Anderson, 2008). Thereby, the three core objectives of information security are confidentiality, integrity and availability (Cawthra et al., 2020). Confidentiality guarantees that secret information is only accessible to intended parties. Integrity refers to the dependability of information and prevents undesired modification of trusted data. Finally, availability ensures that information and information systems are accessible and operational, preventing disruptions that could result in the unavailability of critical resources. Without the goal of availability, security could be simply achieved by preventing access to sensitive information for any entity. Other security objectives can be derived from the three core objectives: Privacy ensures the confidentiality of private data (Spiekermann, 2012), authenticity ensures the integrity of a message’s content and origin, Accountability ensures the availability and integrity of an operation’s actor and non-repudiability ensures the accountability of a message’s sender or receive (International Organization for Standardization and International Electrotechnical Commission (IEC)(2022), ISO).
2.2.1. Requirements & Threat Modeling
Before analyzing, enhancing, or building a system, the security requirements are determined and inferred from higher-level goals (Türpe, 2017). For example, laws might demand privacy measures that necessitate controls at the enterprise level which again lead to requirements at the software level (Deng et al., 2011).
Requirements can also be identified after examining potential threats to security. Threat modeling (Shostack, 2008) is a fundamental methodology to comprehensively evaluate the security of a system by envisioning potential attack vectors and identifying gaps in the defense. Threat modeling paired with the computation of probabilities enables the assessment of security risks. Attack patterns can also serve for detection during runtime.
After analyzing and planning security measures for a complex system, usually multiple defense mechanism are required to address all potential threats, as the options for an attacker are manifold and can target different system layers and components. A well-defined countermeasure is therefore often accompanied by an attacker model (Rocchetto and Tippenhauer, 2016) that defines the capabilities and limitations of potential attackers it aims to defend against.
2.2.2. Access & Information Flow Control
A fundamental concept to achieve basic information security goals is access control (Sandhu and Samarati, 1994), which limits the access to information for certain entities. The formal foundations can be traced back to the Bell-LaPadula model (Bell and LaPadula, 1973) (confidentiality) and the Biba model (Biba, 1977) (integrity). Access control systems regulate the physical or digital access to places and resources with rules expressed in policies. The access policies can be based on identities, roles, or attributes. Multi-user operating systems, for example, restrict the access to data for users and differentiate between permissions for read, write, and execute.
Similar to access control, but acting at the level of program variables is information flow control (Sabelfeld and Myers, 2003). In particular with indirect dependencies, the flow of sensitive information is difficult to track and protect in complex software systems and errors can lead to violations of security policies. Beginning with the seminal work of Denning and Denning (Denning and Denning, 1977), information flow control tracks the propagation of information in a program and verifies its compliance with policies. These policies ensure either integrity by prohibiting untrusted input to influence trusted output, or confidentiality by enforcing the public output to be independent of secure input.
2.2.3. Cryptography
Eventually, secure systems have to transmit sensitive information over insecure channels, e.g., the internet or store it on unprotected devices. Cryptography with primitives like symmetric or asymmetric ciphers or hash functions allows maintaining security in these circumstances. For example, encryption ensures the confidentiality of private data by making it indistinguishable from random data. With digital signatures (Cheon et al., 2023), the integrity of messages can be verified. A threat to cryptographic implementations are side-channel attacks which use unintended information channels of software and hardware systems to extract sensitive information (Krausz et al., 2023). Well known examples are attacks utilizing timing side-channels (Kocher, 1996) or power side-channels (Kocher et al., 1999). When attacking an implementation using a side-channel, the adversary does not attempt to break the cryptographic function itself but exploits the weaknesses of its implementation and the unintended leakage of information produced by it. For example, differences in execution time of a cryptographic function that are secret dependent, can be exploited to infer the secret itself.
Advanced cryptography includes concepts like secure Multi-Party Computation (MPC) (Goldreich, 1998) and Fully Homomorphic Encryption (FHE) (Gentry, 2009). MPC enables a group of mutually distrustful parties to compute an arbitrary, joint function of their private data. The data is contributed by each party without revealing it to any other party, not even to a trusted third party. FHE even allows arbitrary computations on encrypted data without decryption, thus preserving confidentiality.
2.2.4. Intrusion Detection
Besides preventive measures, security can be enhanced with detection methods. An intrusion detection system tries to identify attacks at runtime. The most prominent examples are network intrusion detection systems observing network packages and host-based intrusion detection systems, which observe data files. Intrusion detection can be classified (Axelsson, 2000) into anomaly and signature (pattern) based approaches.
3. Methodology
We used a systematic review methodology (Ralph, 2021) to investigate the state of the art in security DSLs. In what follows, we first describe our DSL selection criteria and the corresponding selection process, followed by a description of the analysis methodology of the selected DSLs.
3.1. DSL Selection
For our systematic review, we searched for security DSLs that can be used by developers, security professionals, and other stakeholders in the development of secure software-intensive systems. We focused on DSLs that explicitly address the security of the system. To this end, we considered all security DSLs that met the following two inclusion criteria
I1.:
The DSL will be used in the development of a secure software-intensive system.
I2.:
The DSL explicitly expresses one or more security aspects.
Figure 1. Literature filtering process.
To systematically identify relevant security DSLs for in-depth analysis, we searched the academic literature on security DSLs. In order to obtain a broad initial coverage of the landscape of security DSLs, we decided to use a generic search term in the initial search and then filter out irrelevant results. Based on the initial dataset of publications identified in this way, we performed an iteration of backward snowballing (Wohlin, 2014) to ensure that we did not miss relevant publications on security DSLs. Figure 1 visualizes the individual sampling and filtering steps performed. In the first step, the initial search, we used the following search term on Semantic Scholar:
security domain-specific language
To avoid dated DSLs that had only low-impact, assure the quality of the papers, and avoid papers from other domains or not focusing on usable security DSLs as aimed by us according to the inclusion criteria, we defined six exclusion criteria.
E1 – published before 2010.:
In our initial search, we excluded papers published before the year 2010. However, we did not apply this exclusion criterion in the snowballing, therefore still including relevant older publications in our data set.
E2 – not in English.:
We excluded papers not written in English for practical reasons.
E3 – low quality.:
To ensure a certain level of scientific quality, we excluded papers that were not peer-reviewed or published in a predatory venue or by a predatory publisher. Please note that the security DSLs themselves must not be the product of scientific papers but must have been discussed in at least one scientific paper.
E4 – not computer science.:
Since we focus on security DSLs in software development, we excluded papers that do not belong to the field of computer science.
E5 – no security DSL.:
We excluded all papers that are not explicitly about security DSLs but only mention these on the side, focus only on languages or only on security but not DSLs.
E6 – no usable DSL.:
Since we aim at practically applicable security DSLs, we excluded papers that do not describe a usable DSL but only theoretical concepts or models.
We searched Semantic Scholar with a python script using the public API of Semantic Scholar and received 14,255 documents. According to exclusion criteria E1, E2 and E4, we limited this search to papers published in 2010 or later, papers in English, and the field of computer science. Based on this accumulation we proceeded with automated filtering, at first removing duplicates, which reduced the number of papers to 811. From an initial manual inspection, we gathered a set of common keywords that indicate research irrelevant for this survey according to exclusion criterion E5:
natural language, language model, language processing, security pattern,
machine translation, linguistic, speaker recognition, english language
We then automatically scanned the title and abstract and paper for these keywords and reduced the papers to 670. From these results we continued with a manual selection based on the title and abstract. We omitted remaining papers not written in English (E2). Furthermore, we looked for indications that the paper either presents a security DSL, discusses the usability or integration of an existing DSLs or surveys DSLs in a security domain and excluded all papers not meeting on of these three (E5 and E6). With 133 papers left we narrowed down the set with a manual selection based on the full paper. In doing so, we also discarded papers, that were not published at a scientific conference or journal, in particular, we discarded three papers, which were solely published as preprints. We rejected three papers, which were published at predatory journals according to Beall’s list [^1] Beall’s list of predatory publishers: https://beallslist.net/ (E3). Multiple publications were excluded, simply because they do not cover a security DSL (E5), although the title and abstract gave indications for this. We excluded several papers, that presented type systems or formal calculi without a usable language (E6). We also dismissed languages and tools that, although used in a security context, come without security specific language features, but for example only a backend which provides security features. Prepose (Figueiredo et al., 2016) is an example for this, as the system has clearly a security goal, but the DSL is completely security agnostic. For three papers, we were not able to access the full document and therefore omitted them.
After this filtering of the initial search results, we then added papers that were referenced as related work in our selected papers in one iteration of backwards snowballing. This step brought up 169 new papers, for which we applied the same manual filtering, leading to 95 additional relevant papers. We ended up with 148 relevant papers covering 120 unique DSLs.
3.2. DSL Analysis
For the analysis of the identified security DSLs, we followed a four-step procedure.
1. Definition of aspects: For each research question, we defined relevant aspects according to which we analyzed the identified security DSLs. We outlined these aspects already when discussing the research questions in Section 1 and provide detailed reasoning for each of them in Section 4.
2. Free text responses: To describe each DSL according to these aspects, we first searched the corresponding papers from our dataset for relevant information to describe the DSL according to each aspect. If we did not find the information in the paper, or if the aspect was open-ended, e.g., whether there are practical applications of the DSL, we searched for further publications on the DSL and for websites or documentation of the DSL.
In this first step, we collected short free-text responses to each aspect. To do this, we divided the selected papers between the first two authors of this paper. While collecting the information on the security DSLs, the two authors were continuously synchronizing to discuss interpretations of single aspects for specific DSLs and to avoid divergence in the collected information.
3. Categorization: After collecting information for each security DSL, we discussed all captured free texts and derived categories from them. Since not all captured information was easily mappable to the identified categories and to improve data quality, the categorization was followed by a data sanitation step in which the first two authors this time divided the individual aspects between them and went through all captures free text answers as well as the full papers again to assign every DSL, depending on the aspect, one or multiple of the identified categories. Again, we supported this procedure with continuous discussions between the first and second author.
4. Statistical analysis: Finally, after classifying all DSLs according to the categories identified for each aspect, we statistically analyzed the captured data. To do this, we created statistics on the distribution of each category for each aspect but also calculated correlations among the different aspects. We present these outcomes in detail in Section 4 and discuss implications in Section 5.
4. Systematic Literature Review
In our SLR, we identified 120 security DSLs which we analyzed in depth to answer the RQs presented in Section 1. In what follows, we present the data extracted by this analysis.
4.1. RQ1: What Security DSLs have been presented in the scientific literature?
Table 1. Security DSLs identified in the SLR (Part 1/2). Categories: Access Control (AC), Cryptography (CR), Information Flow Control (IFC), Intrusion Detection (ID), Requirements (R), Threat Modeling (TM); SDLC Phases: Requirements (R), Design (D), Code (C), Test (T), Deployment (DP), Runtime (RT)
DSL / Author Year Summary Cat. SDLC References Repo.
ASL 1997 generic modeling of access control policies AC D, C (Jajodia et al., 1997)
Bro (Zeek) 1999 detection of network attackers by passive monitoring ID C, RT (Paxson, 1999) (Sharma et al., 2024)
AdeLe 2001 model a database of known attack scenarios TM, ID C, RT (Michel and Mé, 2001)
Ponder 2001 generic modeling of access control policies AC D, C (Damianou et al., 2001)
SPL 2001 single description of global security policies AC D, C (Ribeiro et al., 2001)
CORAS 2002 risk assessment based on semiformal system model TM R, D (Lund et al., 2011; Fredriksen et al., 2002) (SINTEF, 2014)
SecureUML 2002 modeling access control policies and their integration AC D, C, RT (Basin et al., 2011; Lodderstedt et al., 2002; Basin et al., 2006, 2007) (Doser, 2007)
STATL 2002 intrusion detection of known attack scenarios ID C, RT (Eckmann et al., 2002)
UMLsec 2002 planning of secure system design R D (Jürjens, 2002, 2001) (Jürjens et al., 2024)
FlowCAML 2003 secure information flow language IFC C (Pottier and Simonet, 2003) (Simonet, 2003)
uniframe 2003 security contracts for component descriptions AC D (Burt et al., 2003)
Ray et al. 2003 composition of access control frameworks AC D (Ray et al., 2003)
Rei 2003 policies applicable to pervasive environments AC C, RT (Kagal et al., 2003; Kagal, 2002)
XACML 2003 common language for expressing security policies AC C, RT (Ferdous and Poet, 2014; Lorch et al., 2003; Parducci and Lockhart, 2017) (OASIS, 2021)
ESRML 2004 enterprise information security requirements for IS0-17799 R, AC R (Roy et al., 2004)
SAML 2005 exchanging authentication and authorization data AC C, RT (Ferdous and Poet, 2014; Hughes et al., 2005) (OASIS, 2005)
X-GTRBAC 2005 expressive specification of RBAC AC D (Bhatti et al., 2006, 2005a, 2005b, 2008)
WS-CoL 2005 policy assertions of user requirements for web services R C, RT (Baresi et al., 2005)
UMLintr 2006 specification of intrusion detection ID R, D, C, RT (Hussein and Zulkernine, 2006)
WS-SecurityPolicy 2006 security constraints and requirements for web services R C, RT (Bajaj et al., 2006; Nadalin et al., 2009) (OASIS, 2012)
AsmLSec 2007 specifying attack scenarios ID T (Raihan and Zulkernine, 2007)
OSL 2007 specify usage control policies AC D, C (Hilty et al., 2007)
SECTET 2007 policies for secure interorganizational workflows R, AC R, D (Haley et al., 2006; Alam et al., 2006, 2007b, 2007a; Breu et al., 2007)
SecureTROPOS 2007 identification and planning of security requirements R R, D (Mouratidis et al., 2005; Mouratidis and Giorgini, 2007; Matulevicius et al., 2012) (Mouratidis, 2015)
SMCL 2007 secure multi party computation CR C (Nielsen and Schwartzbach, 2007)
Rodríguez et al. 2007 embed security requirements into business processes R R, D (Rodríguez et al., 2007)
WS-Policy4MASC 2007 specification of monitoring and control policies R D, C (Tosic et al., 2007)
CA-RBAC 2008 context-based access control AC C, RT (Kulkarni and Tripathi, 2008)
SELinks/FABLE 2008 enforcement of security policies in web applications AC, IFC C (Afshar and Fallah, 2011; Swamy et al., 2008) (Corcoran et al., 2008)
SFDL 2.0 2008 secure multi party computation CR C, RT (Ben-David et al., 2008; Malkhi et al., 2004) (Tang et al., 2015)
Breu et al. 2008 guide security investment decisions TM R, D (Breu et al., 2008)
USE 2008 specify and verify security policies AC D, C, RT (Sohr et al., 2008)
Cryptol 2009 functional correctness of cryptography CR C (Erkök and Matthews, 2009) (Galois Inc., 2024)
Fabric 2009 secure information flow in distributed systems IFC C (Liu et al., 2009, 2017) (Myers et al., 2019)
Nego-UCONABC 2009 access control for cloud services AC D (Chen et al., 2009)
ReAlSec 2009 identify countermeasures for known attack patterns TM T, DP (Hamdi et al., 2009)
S-Promela 2009 generic security policies AC C (Abbassi and El Fatmi, 2009)
Security Policy Model 2009 generating security policies from security requirements R, AC R, D, C, RT (Wolter et al., 2008; Menzel et al., 2009; Wolter et al., 2009)
SPDL 2009 simplifying the specification of SELinux security policies AC C, RT (Nakamura et al., 2009) (Ynakam, 2008)
Pavlich-Mariscal et al. 2010 access control policies AC D, C (Pavlich-Mariscal et al., 2010)
ADTrees 2010 specify and analyze attack-defense scenarios TM D (Kordy et al., 2010)
ASLan++ 2010 security goals of distributed system designs AC D (von Oheimb and Mödersheim, 2010)
CapDL 2010 capability-based access control AC D, RT (Kuz et al., 2010) (McLeod et al., 2024)
CRePE 2010 fine-grained security policies for android applications AC C, RT (Conti et al., 2010)
SecPAL 2010 simple but expressive, declarative authorization language AC C (Becker et al., 2010)
SSMF 2010 identification of possible attack vectors TM R (Li et al., 2010)
TASTYL 2010 secure multi party computation CR C, RT (Henecka et al., 2010)
X-Policy 2010 dynamic execution permissions for collaborative systems AC D (Qunoo and Ryan, 2010)
Caisson 2011 secure information flow language IFC C (Li et al., 2011)
cPLC 2011 generate secure code for cryptographic two-party protocols CR D, C (Bangerter et al., 2011)
Asnar et al. 2011 identification of permission-related threats R, TM R (Asnar et al., 2011)
UACML 2011 access control modeling language independent of model AC D (Slimani et al., 2011)
Bain et al. 2011 high-level code for secure execution platforms (MPC & FHE) CR C (Bain et al., 2011)
VALID 2011 security goals for virtualized environments R DP, RT (Bleikertz and Groß, 2011)
PTaCL 2012 security policies based on attribute-based access control AC D (Crampton and Morisset, 2012)
Hoisl et al. 2012 secure object flow in business process IFC D (Hoisl et al., 2014)
STRoBAC 2012 spatial temporal role-based access control AC D, C, RT (Thi et al., 2012)
TTCN-3 2012 generation of penetration tests with attack patterns TM T (Stepien et al., 2012) (ETSI, 2023)
Mitchell et al. 2012 secure execution platform based on MPC or FHE CR C (Mitchell et al., 2012)
BCDSDL 2013 network availability with respect to botnet attacks IDE D (Huan et al., 2013)
context-aware RBAC 2013 context-aware access control policies AC D (Schefer-Wenzl and Strembeck, 2013)
CIL 2013 specification of access control policies for Android AC C, RT (Bugiel et al., 2013) (SELinux Project, 2023)
CySeMoL 2013 probabilistic vulnerability assessment TM D, C (Holm et al., 2015; Sommestad et al., 2013) (KTH Network and Systems Engineering Group, 2020)
DJS 2013 solution for untrusted code in javascript CR C (Bhargavan et al., 2013)
FAL 2013 secure logging configurations CR C, RT (Zawoad et al., 2013)
Paragon 2013 practical information flow control IFC C (Broberg et al., 2013a) (Broberg et al., 2013b)
Table 2. Security DSLs identified in the SLR (Part 2/2). Categories: Access Control (AC), Cryptography (CR), Information Flow Control (IFC), Intrusion Detection (ID), Requirements (R), Threat Modeling (TM); SDLC Phases: Requirements (R), Design (D), Code (C), Test (T), Deployment (DP), Runtime (RT)
DSL / Author Year Summary Cat. SDLC References Repo.
RBAC DSL 2013 specification of RBAC policies AC D, C (Kuhlmann et al., 2013)
SI* Asset & Trust 2013 identify insider threats TM R (Paci et al., 2013)
Adão et al. 2013 secure key management CR C (Adão et al., 2013)
XFPM-RBAC 2013 location-aware RABAC AC D (Unal and Çaglayan, 2013)
ZQL 2013 privacy-preserving data processing CR C (Fournet et al., 2013)
BreakGlassPolicies 2014 break-glass policies in role-based access control AC R, D (Schefer-Wenzl and Strembeck, 2014a)
DelegationRBAC 2014 definition of rights delegations AC R, D (Schefer-Wenzl and Strembeck, 2014b)
DPSL 2014 facilitate specification of SOA security policies R D, RT (Wu-Lee and Hwang, 2014)
Haka 2014 describe protocols and apply security policies on network traffic ID C, RT (Arkoon Network Security and ParisTech, 2014) (Fariello et al., 2016)
Jif 2014 practical information flow control IFC C (Pullicino, 2014) (Myers et al., 2022)
Sapper 2014 timing-sensitive information flow control in hardware IFC C (Li et al., 2014)
SecDSVL 2014 specification of security requirements and threat modeling R, TM R, D (Almorsy and Grundy, 2014) (Almorsy, 2013)
Wysteria 2014 secure multi party computation CR C, RT (Rastogi et al., 2014) (Rastogi et al., 2017)
XML Security DSL 2014 simplifying the use of XML security mechanisms CR C (Schefer-Wenzl and Feiertag, 2014)
AIDD 2015 rule-based attack description and response language ID, TM C (Souissi, 2015)
KEYRASE 2015 secure information erasure in software IFC, C C (Askarov et al., 2015)
SecPDL 2015 distributed security policy self-management AC D (Ma et al., 2015)
PPLv2 2015 access and usage control AC C (Cerbo et al., 2015; Trabelsi et al., 2011)
SecVerilog 2015 timing-sensitive information flow control in hardware IFC C (Zhang et al., 2015) (Ferraiuolo, 2015)
STS 2015 modeling and reasoning about security requirements R R, D (Paja et al., 2015) (Giorgini et al., 2020)
SysML-Sec 2015 security requirements and identification of possible attacks R, TM R, D, C (Roudier and Apvrille, 2015)
GemRBAC-DSL 2016 expressive specification of RBAC AC D, RT (Fadhel et al., 2016) (Fadhel, 2017)
Lifty 2016 automated repair for information flow control IFC C, RT (Polikarpova et al., 2020)
LJGS 2016 static information flow control supported by dynamic checks IFC C (Fennell and Thiemann, 2016a) (Fennell and Thiemann, 2016b)
pwnPr3d 2016 attack graph generation TM D (Johnson et al., 2016) (Pwn, 2017)
IoTsec 2017 specification of security requirements R R, D (Robles-Ramirez et al., 2017)
Jasmin 2017 efficient, safe, secure and correct cryptographic implementation CR C (Almeida et al., 2017) (Laporte et al., 2024)
Low* 2017 efficient, safe, secure and correct cryptographic implementation CR C (Protzenko et al., 2017) (Martínez et al., 2024)
ProScript 2017 secure cryptographic protocol code CR C (Kobeissi et al., 2017) (Kobeissi et al., 2022)
Vale 2017 secure and efficient cryptographic implementation CR C (Bond et al., 2017) (Baumann et al., 2024)
SRESOFL 2017 formal specification of security requirements R R, D, C (Emeka and Liu, 2017)
STREAMS 2017 specify information flow control policies for hardware IFC C (Kwon et al., 2017)
ChiselFlow 2018 timing-sensitive information flow control in hardware IFC C (Ferraiuolo et al., 2018)
IoT Language 2018 threat modeling and risk assessment TM D (Beaulaton et al., 2018) (Beaulaton, 2018)
LOCKS 2018 formal and convenient security goal specification R, TM R, D, C (Kumar et al., 2018)
LUCON 2018 data-centric security policies for distributed systems IFC C (Schütte and Brost, 2018)
MAL 2018 threat modeling and attack simulation TM D, T (Katsikeas et al., 2019; Johnson et al., 2018; Widel et al., 2023) (Wällstedt et al., 2024)
FaCT 2019 timing side-channel secure cryptography CR C (Cauligi et al., 2017, 2019) (Soeller et al., 2019)
Flamio 2019 coarse-grained, dynamic, information-flow control IFC C (Pedersen and Chong, 2019) (fla, 2018)
IDS-DL 2019 visual description of intrusion detection systems ID D (Amine et al., 2019)
UML-SR 2019 specification of low-level security requirements R R, D (Mohsin and Khan, 2019)
Cao et al. 2019 access control between system components AC D (Cao et al., 2019)
Zhao et al. 2019 integration of functional and security models R D (Zhao et al., 2019)
QIF-RTL 2019 quantitative information flow control in hardware IFC C (Guo et al., 2019; Reimann et al., 2021) (Guo and Jin, 2019)
Zee 2019 information flow control for software systems with RTE IFC C (Vorreiter Pedersen and Askarov, 2019) (zee, 2019)
DynaMo 2020 policy analysis and automated source code generation AC C (Amthor and Schlegel, 2020)
Legal-GRL 2020 legal compliance requirements R, AC R, D, C, T (Sartoli et al., 2020)
MACasync 2020 secure information flow with respect to concurrency IFC C (Tom’e Corti nas et al., 2020)
REAP 2020 specify structure of security policies AC D (Amthor and Schlegel, 2020)
Co-Inflow 2021 dynamic, coarse-grained information flow control IFC C (Xiang and Chong, 2021) (Xiang, 2021)
JE 2021 language support for TEE integration IFC C (Oak et al., 2021) (Oak et al., 2022)
SSEL 2021 specification of secure SoC component communication R R, D, C, T (Raj et al., 2021)
Hamraz 2022 information flow control in distributed system IFC C (Li et al., 2022)
IFL (IFCIL) 2022 information flow control for SELinux IFC C (Ceragioli et al., 2022)
Before examining the identified security DSLs in detail, to get an understanding of the background of the examined security DSLs, we provide an overview of the DSLs, the scientific communities from which they originate, and how many of them are still actively maintained and accessible.
Summary of the identified DSLs
Tables 1 and 2 show the security DSLs identified in the systematic literature review as well as the corresponding references. While DSLs, such as SELinks/FABLE, SecureUML, or CORAS, were presented in multiple papers that we found in our literature review, two papers contain two DSLs in a single paper (DynaMo & REAP (Amthor and Schlegel, 2020) and XACML & SAML (Ferdous and Poet, 2014)). The abbreviation of the Security Policy Description Language (Ma et al., 2015) conflicts with the abbreviation of the Simplified Policy Description Language (Nakamura et al., 2009), we therefore refer to the former with SecPDL and only the latter with SPDL. As visualized in Figure 3, the DSLs identified in our literature review were published between 1997 and 2022. It is clearly visible that the number of published DSLs increases relatively steadily until a peak in the publication of security DSLs is reached in 2013. Thereafter, the number of newly published security DSLs decreases steadily. This raises the hypothesis that security DSLs have been published for most relevant security aspects. We will explore this hypothesis in the detailed analysis of the security DSLs.
Background of the DSLs
Most papers (49 %) have been published on venues targeting the security community, followed by 11 % of venues in software engineering. Networks (9 %), applied computing (7 %), information systems (7 %), programming languages (7 %), and distributed computing (6 %) venues are other commonly observed communities. Note that a single venue may target multiple communities. A significant portion of the venues (3 %), we could not assign to specific communities because they are general in nature and do not target identifiable communities, but rather computer scientists in general. For example, this journal falls into the category of publication venues targeting computer scientists in general. In Figure 2(b), we list all conferences and journals where at least four of the papers we examined were published, as well as the communities to which we assigned them.
0010101010202020203030303040404040505050506060606070707070SecuritySoftware EngineeringNetworkProgramming LanguagesInformation SystemsApplied ComputingDistributed ComputingTheoretical Computer ScienceOperating SystemsHardwareRequirementsArtificial IntelligenceOtherNumber of Papers
(a) Communities in which was published
Venue Name Community Papers
Computer Security Foundations Symposium (CSF) Security 8
Conference on Availability, Reliability and Security (ARES) Security 6
Symposium on Security and Privacy (S&P) Security 5
USENIX Security Symposium Security 5
Conference on Computer and Communications Security (CCS) Security 4
Symposium on Policies for Distributed Systems and Networks (POLICY) Distributed Computing, Network, Security 4
(b) Venues with four or more papers
Figure 2. Communities of the publications from which the DSLs originate
Only 40 of the 120 identified security DSLs (33 %) have a publicly accessible repository or website. Note that our initial search for publications covers the time span between 2010 and 2022 and with the backwards snowballing we additionally identified relevant papers dating back until 1997. Many security DSLs have been updated even after their release, as shown in Fig. 3. In particular, 25 % of the updates happened in 2023 or the first half of 2024, indicating that quite a few of them are actively maintained and possibly also used in practice or academia.
19971997199719971998199819981998199919991999199920002000200020002001200120012001200220022002200220032003200320032004200420042004200520052005200520062006200620062007200720072007200820082008200820092009200920092010201020102010201120112011201120122012201220122013201320132013201420142014201420152015201520152016201620162016201720172017201720182018201820182019201920192019202020202020202020212021202120212022202220222022202320232023202320242024202420240022224444666688881010101012121212Number of DSLspublishedlast update
Figure 3. Publication year of the DSLs and latest update on their publicly available repository or website
Observation 1. Security DSLs were primarily published in the mid-2010s, mainly from the security community, but also from other communities such as software engineering. While few DSLs are publicly available, many of those that are are still actively maintained.
4.2. RQ2: What security aspects are targeted by the security DSLs?
In this work we only cover DSLs that are security specific. To get an overview over the security aspects covered by the DSLs we analyze their fundamental security objective and provide a coarse-grained categorization according to the security concept and the system scope targeted by the DSLs. Furthermore, we capture the availability of attacker models and defense mechanisms.
First, we distinguished between problem and solution oriented DSLs. Most languages (74 %) focus on solutions, e.g., by enhancing defense mechanism implementation and 23 % of the DSLs focus on (identifying) problems. A few DSLs (3 %) are both, problem and solution oriented.
Security objectives
We investigated which security objectives the security DSLs address and were able to map the DSLs to seven fundamental security objectives (cf. Fig. 4(a)).
A DSL can address multiple security objectives, some DSLs are not tied to a specific objective and can be used for any potential security objective. This can be the case, for example, for threat modeling languages. However, the security objectives are explicitly formulated for only 46 % of the security DSLs, sometimes by making them part of the security DSL or the framework around. The other 54 % of the DSLs address security objectives implicitly, e.g., in terms of considered countermeasures.
Most DSLs target one of two core security objectives: confidentiality and integrity. The former is addressed by 55 DSLs (46 %) and the latter by 37 DSLs (31 %). Availability, the third core security objective is only targeted by 13 DSLs (11 %) and derived security objectives like privacy are focused by even fewer languages. However, 48 % of the DSLs do not consider a fixed set of objectives but can potentially target any security objective.
For example, confidentiality is assured by languages for information flow control. LABEL:lst:paragon gives instances of illegal direct and indirect (via branching) flow of secret information in the Paragon language. It demonstrates how features like explicit declassification can be used to allow flow from secret to public values, which can be necessary in real-world applications.
Listing 1: An example in the Paragon language that demonstrates legal and illegal flow of data by direct and indirect flow including the feature of declassification (from Broberg et al. (Broberg et al., 2013a)).
1void method(?Declass.high int highData) {
2 Declass.low int lowData;
3 lowData = highData; // Illegal direct flow
4 lowData = Declass.declassify(highData); // OK
5 if (highData == null) {lowData = null;} // Illegal indirect flow
6}
0055551010101015151515202020202525252530303030353535354040404045454545505050505555555560606060ConfidentialityIntegrityAvailabilityAccountabilityAuthenticityNonrepudiabilityPrivacyAnyNumber of DSLs
(a) Fundamental security objectives
0055551010101015151515202020202525252530303030353535354040404045454545RequirementsAccess ControlThreat ModelingIntrusion DetectionInformation Flow ControlCryptographyNumber of DSLs
(b) Categories of security concerns
Figure 4. Security aspects addressed by the security DSLs
Categories of security concerns
Furthermore, we investigated which security concerns the different DSLs address. We identified six different categories of targeted security concerns (cf. Fig. 4(b)). Thereby, a single security DSL can target multiple of the identified categories.
Requirements.:
Of the 120 DSLs we discovered, 23 focus on security requirements. Some of them are tied to business process models, while others standardize the specification of requirements with policies, for example for web services. A typical example for this category is SecureTROPOS (Mouratidis and Giorgini, 2007), a visual, graph-based DSL aimed at the identification and validation of security requirements for multiagent systems.
Access control.:
Access control is the most dominant category we identified (43 DSLs). Multiple DSLs of this category focus on effective specification of access control policies. Furthermore, languages in this category are used to define access control models, such as the UML-based access control modeling language (UACML(Slimani et al., 2011).
Threat modeling.:
We found 17 DSLs that can be placed in the threat modeling category. Some of them facilitate the identification of attack vectors based on design models. Others allow efficient descriptions of attacks for simulations and to detect potential countermeasures. Part of this category is, e.g., MAL (Johnson et al., 2018), a probabilistic meta threat modeling and attack simulation language
Intrusion detection.:
DSLs for intrusion detection systems are used to specify known attack patterns which can then be identified during runtime by the intrusion detection system and yield a response. For example, Listing LABEL:lst:statl demonstrates how the fundamental concepts of states and their transitions in STATL (Eckmann et al., 2002) allow expressing potentially vulnerable network protocol states to detect attacks. We mapped 9 DSLs to this category.
Information flow control.:
Languages for information flow control are typically equipped with a type system that features dedicated types to address confidentiality (secret vs. public) and integrity (trusted vs. untrusted). Via type inference, the flow of sensitive information over variables is automatically tracked, and the framework ensures dynamically or statically that, e.g., secret variables do not reach publicly observable outputs. We mapped 23 DSLs to this category; a well known representative is Jif (Pullicino, 2014).
Cryptography.:
Cryptographic primitives are usually described with a mixture of continuous text, pseudocode and mathematical expressions, leading to ambiguous specifications that are not suitable to formally verify the equivalence of their implementations. This category includes DSLs tackling this problem by including language constructs that enable an explicit and efficient expression of the underlying primitives (Cryptol (Erkök and Matthews, 2009)). LABEL:lst:cryptol shows how a part of the AES encryption can be specified with Cryptol using its built-in type for polynomials, primitive polynomial modulus, and multiplication operations. Other DSLs in this category enable a simple and correct usage of cryptography, or focus on side-channel secure implementations of cryptography (FaCT (Cauligi et al., 2019)). Some DSLs allow automated generation of MPC or FHE protocols from the function that is computed on the sensitive data (Wysteria (Rastogi et al., 2014)) and thus allows non-cryptographic experts to use this advanced cryptography. In total, 19 DSLs focus on cryptography.
Listing 2: State transition in STATL to identify half-open TCP connections (from Eckmann et al. (Eckmann et al., 2002)).
1transition RST (s1 -> s0) unwinding {
2 [IP ip [TCP tcp]] :
3 (ip.header.src==victim_addr) &&
4 (tcp.header.src==victim_port) &&
5 (ip.header.dst==attacker_addr) &&
6 (tcp.header.dst==attacker_port) &&
7 (tcp.header.flags & TH_RST) }
Listing 3: Specification of multiplication in the Galois Field including the irreducible polynomial for AES written in Cryptol (from Cryptol (Erkök and Matthews, 2009))
1 poly = <|x^^8+x^^4+x^^3+x+1|>
2 gf28Mult : (GF28, GF28) -> GF28
3 gf28Mult (x, y) = pmod (pmult x y) poly
Our categorization covers a wide range of subdomains within information security and indicates that DSLs have already been proposed and developed for most relevant security aspects. It is noteworthy that most categories discussed above represent some type of defense mechanism (Access Control, Intrusion Detection, Information Flow Control, and Cryptography), thus potentially indicating that security DSLs may focus on concrete defense mechanisms. In fact, 82 % of the DSLs support a concrete defense mechanism. Many languages that are not associated with a specific defense mechanism aim to specify security requirements, and the choice of defense mechanisms is left to the user. Nevertheless, defenses play an important role in the use of these DSLs as well.
System Scope
Security concerns and defenses are applied with different scopes of the information system. For example, threat modeling can be performed using the same principles at a high level, where the entire enterprise is in scope, and at a lower level, where only the software is considered. Different system scopes present unique challenges, for example, quantifying probabilities for human behavior is likely to be more difficult than quantifying probabilities of success for brute force attacks. We used a coarse-grained categorization of information system scopes that reflects the common structuring of security research domains. We mapped each DSL to a single scope that most closely matches the target scope of its security concept or countermeasure.
Enterprise.:
At the highest level, with 19 languages (16 %), the entire enterprise is in scope of the security concept of the DSL. Relevant aspects include business processes and human interactions.
IT System.:
We have sorted 73 DSLs (61 %) into the Information Technology (IT) system scope. In this category multiple subsystems combined over, e.g., a network are relevant, but also the interaction of multiple programs, including operating systems.
Software.:
Twenty DSLs (17 %) focus on the security of individual software, for example, software implementations of cryptographic primitives.
Hardware.:
Although this work aims at DSLs for software-intensive systems, we also included 8 (7 %) languages that focus on the security of individual hardware components. These are primarily (programming-language like) hardware description languages.
In summary, most DSLs facilitate security with a scope that considers multiple interacting components of the system. This is intuitively the case for DSLs supporting the requirements and design phases of secure information systems, but can also apply to programming languages with security mechanisms that have a scope beyond the software.
Attacker model
Ideally, a defense mechanism is tied to an attacker model that defines the kind of attackers it aims to protect against. This includes limitations of the attackers capabilities and allows determining if a countermeasure is sufficient for the applied scenario.
Available.:
Many security DSLs (36 %) come with an explicitly specified attacker model. The categories cryptography and information flow control stand out with 87 % and 84 %, respectively.
Instance-specific.:
For a number of DSLs (15 %) it is not possible to specify a concrete attacker model since it depends on the concrete subject system to which this DSL is applied. These DSLs are mostly design-time security DSLs.
Focus of the DSL.:
There are also some security DSLs (7 % ) that do not come with an attacker model since it is their purpose to allow security experts to specify such an attacker model, e.g., as part of threat modeling.
None.:
Unfortunately, with 51 security DSLs (43 %), a significant number does not provide an attacker model. While for some security DSLs it is not feasible to specify an attacker model, e.g., DSLs for reasoning about security requirements, these are only a few exceptions of the DSLs not providing an attacker model.
Observation 2. Security DSLs exist for a wide range of security objectives, countermeasures, and system scopes. Unfortunately, a significant share lacks a precise attacker model.
4.3. RQ3: What phases of the development process are supported by the DSLs?
To answer the third research question, we first investigated in which phases of the SDLC the security DSLs are used and what artifacts are expressed in them. Thereafter, we investigated with which other artifacts they are integrated and how.
Relevant phases of the SDLC
The DSLs are used in six different phases of the SDLC, whereby 44 % of the DSLs are relevant in multiple phases. On average, every DSL is used in 1.5 phases. Figure 5 shows how the DSLs distribute across the phases of the SDLC and concrete numbers of how often each phase is relevant for a security DSL, in which phases instances of the DSLs are created, and in which phases these are actually put into use.
(a) Use of the security DSLs across multiple phases
Req.DesignCodeTestDepl.Runtime0020202020404040406060606080808080100100100100Number of DSLsinstantiationrelevantuse
(b) DSL usage in the phases
Figure 5. Relevant SDLC Phases for the Usage of Security DSLs
Requirements.:
First, for 22 DSLs (18 %) requirements engineering is a relevant phase. Usually, the DSLs support to identify and specify security requirements for the system to be developed. For example, Fig. 6 shows how the requirements of a health care system are enriched with security requirements, i.e., confidentiality or integrity requirements on the data communicated, using STS. Requirements specified this way are reused in 82 % of the cases in later phases of the SDLC.
Design.:
After the requirements phase, 49 DSLs (41 %) are instantiated in the design phase. Another six DSLs instantiated in the requirements phase are relevant to the design phase (55 DSLs in total). In the design phase, more-detailed security requirements are specified in relation to the system’s design. Here, a wide range of design models with different levels of abstraction are supported. For example, UMLsec allows to explicitly classify information such as sensitive information in abstract domain models and at the same time allows planning detailed requirements on communication security, e.g., whether a connection needs to be encrypted, on detailed deployment diagrams (see Fig. 8). Design-level security DSLs are mainly used in combination with the requirements phase and concrete code. Consequently, 40 DSLs are put into use in this phase. However, their use is also relevant in the test phase and at runtime.
Code.:
With 76 DSLs, the majority of all DSLs (63 %) addresses the implementation phase on source code. Of these DSLs, 67 DSLs are also instantiated in this phase. Code-only DSLs usually provide the means to implement secure code, e.g., containing security labels for secure data flow analysis, or for securely implementing cryptographic code. While most DSLs are only relevant to the code phase, there is still a significant number that also addresses design-time or runtime, or all three of these phases. We observed no DSL that is used in combination with any other phase except unless it also covers the design-phase. Mostly, concrete security objectives are planned at design-time, implemented or verified in code, thereby putting 54 of the relevant DSLs into active use in the implementation phase, and probably also enforced at runtime.
Test.:
Only 6 DSLs (5 %) are used in the test phase, and all of them are used in further phases. However, the majority of these DSLs (4 DSLs) are instantiated in this phase as part of security tests, accompanying other parts of the DSL instantiated in earlier phases. The test phase is not the main focus of these DSLs, but they do allow, among other things, the generation of test code.
Deployment.:
The fewest DSLs, only 2 DSLs (2 %), are used in the deployment phase. Since DSLs are used in the test phase, they are also used in other phases, such as planning security goals for virtualized environments and actually deploying them securely.
Runtime.:
Finally, 52 DSLs are relevant to the runtime of the system, often however the DSL instance is not actively used but transformed into another format before. Only 26 DSLs (22 %) are immediately used at runtime, meaning that the DSL instance is actively used at runtime. For security DSLs at runtime, we can observe two different kinds of usage. First, security requirements specified at design-time being monitored at runtime, and second, concrete monitoring rules implemented in models or code being executed at runtime.
Figure 6. Partial actor diagram for an electronic health care system in STS.
Artifacts specified using the DSLs
Following the development phases in which the security DSLs are used, we observed a wide range of artifacts that are expressed using the DSLs (cf. Figure 7).
[…]
Figure 7. Kinds of artifacts expressed using the security DSLs
Requirements.:
Although 24 DSLs focus on security requirements, only 8 of them (7 % of all DSLs) are used to specify requirements. All other DSLs extend artifacts used in other phases with security requirements, such as in UMLsec (cf. Fig. 8) that enriches UML models with security requirements on communicated data and the security of communication paths used.
Process models.:
Only 2 DSLs express (business) process models with a visual representation of security considerations. While Rodríguez et al. (Rodríguez et al., 2007) extend the BPMN, the Security Policy Model (Wolter et al., 2008; Menzel et al., 2009; Wolter et al., 2009) defines its own process model notation.
Design models.:
With 15 DSLs (13 % of all DSLs), many DSLs are used to create design models that express a secure design of the planned system. For example, as shown in Fig. 8, UMLsec is used to plan secure deployments of software systems.
Code.:
Corresponding to the SDLC phase of coding, where most security DSLs are used, with 32 DSLs (27 %), code is the second-largest artifact category. DSLs are used to create custom code artifacts, such as STATL (see LABEL:lst:statl) or Cryptol (see LABEL:lst:cryptol), as well as code artifacts in common general purpose languages, such as TASTYL which is based on python. We recall that only 17 % of the DSLs facilitate security with a scope limited to software, for example, by addressing side-channel security or memory-safety. Therefore, many DSLs with code artifacts feature defense mechanisms that consider the interaction of several system components, e.g., distributed systems.
Hardware descriptions.:
Hardware description languages that express security considerations of hardware, typically in the form of logic circuits that can be generated and come with some security guarantees, are provided by 5 DSLs (4 %).
Attack/defense descriptions.:
Attacks and, in the case of some DSLs, appropriate defenses can be described using 11 of the DSLs (9 %). Most of these DSLs use a graph-based representation such as attack trees (Kordy et al., 2010) or probabilistic attack-defense graphs (Katsikeas et al., 2019; Johnson et al., 2018; Widel et al., 2023).
Policy files.:
With 43 DSLs (36 %), policy files are the largest artifact category we encountered. Since policies are a fundamental aspect of security and a component of various defense mechanisms, especially access control systems, which were the most common category of security aspects among all security DSLs (see Section 4.2), this is an expected observation. Within this category, 29 DSLs are used to specify plain policies, while 13 DSLs (11 %) are used to create concrete access control models that go beyond simple policies. In addition, PPLv2 (Cerbo et al., 2015) specifies detailed models for access and usage control, which even goes beyond what can be expressed in the policy files or access control models considered above.
Intrusion detection rules.:
Some form of intrusion detection rules can be defined using 3 of the DSLs (3 %). AIDD (Souissi, 2015) and IDS-DL (Amine et al., 2019) facilitate the specification of intrusion detection rules, which could be seen as special policies or attack and defense descriptions. Haka (Arkoon Network Security and ParisTech, 2014) allows specifying communication protocols in combination with filtering rules to identify and address malware network activities.
Security configuration.:
Lastly, FAL (Zawoad et al., 2013) is used to specify configuration files of a secure logger.
Integration with other artifacts
The investigated DSLs are usually not used in isolation but 62 % of all DSLs are integrated with other artifacts. We consider an DSL as being integrated with an artifact if it has an explicit relation to this artifact, e.g., by extending it or explicitly referring to it.
Requirements.:
First, 4 security DSLs (3 %) enrich requirements expressed in different artifacts with security-related aspects such as relevant threats for concrete requirements.
Design models.:
Similarly, 20 DSLs (17 %) enrich design models with relevant security objectives, threats, and planned security features. For example, in Figure 8, the UMLsec security DSL is used to enrich a UML deployment diagram with information on the criticality of communicated information and requirements on the security of used communication paths.
Attack/defense descriptions.:
Only the LOCKS DSL is integrated with attack descriptions in SAM models for specifying security goals.
Code.:
Most DSLs (33 DSLs, 28 %) are integrated with some form of code. One half of these DSLs is complied into code in some programming language that will be embedded into functional source code and the other half directly extends functional source code.
Policies.:
While 13 DSLs (11 %) are integrated with policies, thereof, only 9 DSLs express themselves policies or access/usage control models. Usually, these are integrated or compatible with a policy standard, such as XACML (6 DSLs).
Runtime components.:
Overall, 10 DSLs (9 %) are integrated with runtime components. Mostly, these DSLs are integrated with components of the system at runtime (8 DSLs), usually meaning that the instance of the DSL is executed at runtime in combination with this runtime component, e.g., compiled functional source code. In addition, 2 DSLs are integrated with intrusion detection systems and build the basis for monitoring a system.
Hardware.:
Only STREAMS (Kwon et al., 2017) integrates with existing hardware description artifacts by providing a hardware module for use in Verilog. All other DSLs that target hardware description specify these in their own languages, but do not integrate with existing specifications.
Figure 8. UML deployment diagram of a health records system annotated with UMLsec for planning security of communication paths based on labeled information communicated over them (oriented on Peldszus (Peldszus, 2022))
Unlike the significant number of DSLs that are used in multiple SDLC phases would make one expect, only 7 DSLs are integrated with multiple artifacts. This is due to the fact that we consider a transformation of a DSL instance into another format not as an actual integration a developer has to deal with. DSLs that are integrated with multiple artifacts are usually design models that are extended or used to express requirements or policies.
Kind of integration
Less than half of the DSLs examined (51 DSLs) are relevant to multiple phases of the SDLC as part of the development of secure systems. We identified five variants in which security DSLs can be integrated into development. It is possible that a single DSL can be integrated into the SDLC using multiple of these variants.
Planning.:
Most of the security DSLs (23 %) serve as languages for planning artifacts. The use of DSLs enables the documentation of security requirements, planned countermeasures, and other security aspects that must be considered or realized in subsequent phases of the SDLC. These serve as requirements for subsequent phases, with the manual usage of artifacts created using the corresponding DSLs representing one of their primary purposes.
Enriching artifacts.:
The second-highest number of DSLs (18 %) is integrated into the SDLC by enriching artifacts that are created as part of the development process anyway, e.g., for planning or documentation purposes, with security considerations.
Detailing.:
A significant proportion of DSLs (11 %) permit the continuous detailing of the security aspects expressed using them over the phases of the SDLC. As the development of the system progresses, the DSL instances are enhanced with further information, particularly in relation to access control policies, security requirements, and threat models.
Reuse.:
Many artifacts created using security DSLs (13 %) are reused in later phases of the SDLC by other security engineering techniques, such as access control policies derived at design time that serve as input to an access control engine at runtime.
Traceability.:
Finally, the Access Control Framework (Pavlich-Mariscal et al., 2010) also allows tracing security requirements across different artifacts created in the SDLC.
None.:
The vast majority of the security DSLs (58 %) has been designed with a single artifact in mind and is used in isolation.
Observation 3. Security DSLs are used in all phases of the software development life cycle (SDLC), but 58 % of the DSLs do not feature any means of integration into the SDLC. There is a significant focus on design-time and code-level DSLs. Particularly, deployment and testing are underrepresented. The most common artifacts of security DSLs are security policies, code, or design models. While security DSLs sometimes are used in multiple SDLC phases, this is primarily the case for security DSLs targeting the requirements engineering or design phase.
4.4. RQ4: What types of DSLs exist, and how are they used?
For embedding security aspects into development processes as discussed above, DSLs can be specified following various language design principles. To get a better understanding of the technical perspective on security DSLs, first, we look at how the security DSLs are defined and what are their characteristics. To this end, we discuss whether the DSLs are an extension of an existing language or are newly designed languages and how instances are represented. Thereafter, we discuss how developers work with the DSLs and how they are supported doing so.
Language design
When examining the artifacts created by the security DSLs, we observed that some of them are defined entirely by syntax, while others are enriched versions of other languages, such as extended UML models. In general, language engineering typically distinguishes between two types of language design, both of which we observed in the security DSLs we studied.
External DSLs.:
Most commonly (66 %), we observed external DSLs that come with custom-tailored parsers or interpreters specifically designed for the language features of the DSL.
Internal DSLs.:
Unlike external DSLs, internal DSLs, which make up the remaining 34 % of DSLs analyzed, enrich an existing language and use its language infrastructure. Among the investigated internal DSLs, the most common host language is the Unified Modeling Language (UML), which is extended by 40 % of all internal security DSLs via so-called UML profiles. In total, 50 % of internal security DSLs extend modeling languages, such as SysML, BPMN, and AADL. The second-largest group of internal DSLs extend programming languages (35 %). Thereby, Java and Haskel are the most extended languages (4 DSLs each), followed by ASML, F*, F#, Javascript, Promela, and Scala with one DSL each. Finally, three DSLs each enrich policy languages (IF, WS-Policy, and XCAML) and requirements languages (GRL and two times SI*). For example, Asnar et al. (Asnar et al., 2011) extend the SI* goal modeling language (Massacci et al., 2010) with threat modeling capabilities.
Relation to existing DSLs
The DSLs analyzed are usually new languages, which in the case of internal DSLs may be related to an underlying modeling or programming language. However, 18 % of them are based on an existing DSL and extend it with additional aspects. Extended security DSLs are mainly DSLs for specifying access control policies (31 %), e.g., WS-CoL and WS-Policy4MASC extend the WS-Policy DSL with additional policy concepts. Secure code DSLs are also being extended, e.g., Fabric and $J_{E}$ both extend Jif with additional concepts. Furthermore, with 35 %, the percentage of threat modeling DSLs which extend another security specific language is remarkably high.
Instance representation
The second characteristic of a language is how instances are represented and made accessible to users. This can be entirely textual, visual, or a combination of the two.
Textual DSLs.:
The majority of DSLs (69 %) are textual languages. The DSLs define a textual syntax for the language, allowing users to specify instances in any text editor. DSLs such as Jif extend the language syntax of Java with additional elements and can be used in a similar way to standard Java source code. STARTL, Paragon, Cryptol, and many others define their own programming language-like syntax. Another style of textual syntax is used by WS-Policy, the DSLs that extend it, and many other policy languages. These often use an XML syntax. Almost all DSLs in the information flow control and cryptography categories are textual languages.
Visual DSLs.:
STS and UMLsec are examples for a visual DSLs. Figure 6 provides a partial example of a social diagram in STS. The visual notation of STS is oriented on SI* goal models, while UMLsec extends the UML. In the back, both models are serialized as XML files, but these are not intended for manual use. Overall, 24 % of the DSLs we examined work on a visual representation.
Textual & Visual DSLs.:
A few DSLs (7 %) allow users to work on both kinds of representations. Depending on the concrete task, they can select the more suitable representation of the instance, i.e., users are often faster in writing instances in the textual representation but are better in understanding and debugging them in presented visually.
Interestingly, 64 % of problem-oriented DSLs are visual, and 11 % are textual and visual.
Editing support
To allow users to work efficiently and effectively with DSLs it is important to provide them with editing support. Editing support ranges from simple batch syntax checking, such as usually provided by compilers, over its embedding into IDEs for live feedback on the syntax, and syntax highlighting to semantic feedback on the expressed aspects or suggestions for auto-completion. Among the DSLs, we identified four degrees of editing support (cf. Fig. 9).
[…]
Figure 9. Statistics on editing support and instance representation and language design (from left to right).
None.:
Most DSLs come with no editing support at all (64 %). Users have to write instances with any editor and are not supported in doing so. Mistakes only get recognizable, in the case of errors during compilation, interpretation, or other forms of DSL usage.
Basic.:
At least some basic editing support, usually comprising syntax highlighting and checking, is provided for 16 % of the DSLs.
Extended.:
A few DSLs (8 %) extend basic editing support with tools that facilitate the specification of language instances. This support usually, includes simple features such as the completion of keywords defined in the language, but does not go as far as considering the expressed semantics.
Sophisticated.:
The highest degree of editing support comprises instance-specific checks of the expressed semantics and sophisticated editing features such as auto-completion and is provided by 13 % of the DSLs. This degree of editing support is mainly implemented via language-engineering frameworks such as xtext that automatically generate editing support features or UML-based languages that are enriched with semantic checks embedded via OCL constraints.
Observation 4. Security DSLs in general are usually textual, but most problem-oriented DSLs have a (partially) visual representation. Editing support is only available for 37 % of the DSLs.
4.5. RQ5: What are the semantics of the DSLs?
To better understand how the security DSLs affect the development of secure software systems, we investigated their semantic usage.
Semantic use of the DSLs
First, we looked at the semantic purpose for which the security DSLs are used in the development of secure software systems (see Fig. 10). A single DSL can be used in multiple semantic ways and on average the DSLs are intended to be used in 1.5 ways.
Figure 10. Semantic usage of the DSLs (one DSL can be used in multiple ways).
Knowledge.:
With 29 DSLs (24 %), a significant number of DSLs serve as a knowledge exchange format between stakeholders for documentation purposes. For example, the security requirements specified in the DSLs are given to the developers, who then implement them.
Analyze.:
The majority of DSLs (52 DSLs, or 43 % of all DSLs) provide support to developers by analyzing the software system as represented by an instance of the DSL for security issues, such as for identifying insecure design decisions such as the lack of encrypted communication of sensitive data that should be addressed by software engineers.
Transform.:
There are 24 DSLs (20 %) that are intended to be the source of an automated transformation into another instance format, i.e. another DSL. Typically, these DSLs allow for a simpler specification of the security aspects expressed than the usually more generic target format and domain-specific knowledge that is hidden by the DSL is added through the transformation. In most cases, access control policies derived at design time are transformed into other formats that can be executed at runtime, e.g., by an access control engine.
Codegen.:
The generation of secure source or byte code in a general purpose programming language that then can be embedded by developers into the implementation of the system is supported by 9 DSLs (8 %). The code generated this way is usually not executable on its own but must be manually embedded into the implementation of the software system.
Compile.:
The compilation of an DSL instance into an executable format in machine code is supported by 20 DSLs (17 %). These DSLs assemble executable programs after compilation, e.g. Jif is an extension of the Java programming language with information flow constraints that compiles to Java bytecode can be executed on the JVM.
Interpret.:
Lastly, 33 DSLs (28 %) are executed by an interpreter. Unlike compiled DSLs, no executable machine code is produced, but the DSL is executed as it is by the interpreter. DSLs used in this way mainly cover security constraint specifications such as access control policies that are enforced by a runtime engine.
Automation
Security DSLs are commonly provided with tooling that automates certain aspects of their semantic usage. Among the DSLs that come with such tooling, the average number of automated semantic usages is 1.2, which is much lower than the intended average of 1.5 usages.
None.:
While the semantic usage of capturing knowledge for the use by humans is the only semantic usage of security DSLs that cannot be automated and only 15 DSLs have solely this purpose, 33 DSLs (28 %) do not come with any automation. Often, some automation is outlined in future work but has never been realized.
Analyze.:
The majority of security DSLs (36 %) come with an automation of their proposed analyses, e.g., for identifying security issues. While 52 DSLs propose analysis as semantic usage, with 43 DSLs most of them actually automate this usage (83 %). Analyzes are mainly hard-coded but are also frequently implemented using query or constraint languages such as the OCL.
Transform.:
All security DSLs that propose a transformation, with the exception of four (83 % of the DSLs that propose transformations), automate this transformation. In total, 20 DSLs (17 %) automate transformations. Transformations are mostly implemented in general purpose languages and model transformation languages build the exception.
Codegen.:
All 9 DSLs that propose code generation actually implement tooling for generating source code from the DSL instances.
Compile.:
Only 9 of the 20 of the DSLs that propose to compile the instances into an executable format actually implement such a compiler (60 %).
Interpret.:
Similar to compilers also for only 21 of the DSLs (64 %) for which an interpretation at runtime is proposed such an interpreter is actually implemented.
When it comes to the actual analysis of security DSLs for detecting vulnerabilities or integrating them with other SDLC phases through the transformation into another format, 83 % of the DSLs come with such tooling. However, putting the DSLs actually into action remains mainly on a theoretical basis and is often not implemented (only for 60 % of compilation and 64 % of interpretation). One exception is code generation for usage by developers as part of their implementations which is implemented for all 9 DSLs for which such a code generation has been proposed.
Observation 5. The common semantic purposes of security DSLs are analysis, knowledge exchange, or (either direct or via indirect means like transformations or code generation) code execution. Automation, in particular by compilation and interpretation is often not implemented.
4.6. RQ6: How effective are the DSLs?
In our last research question, we focus on the effectiveness of the security DSLs. First we look at how the effectiveness of the security DSLs is assessed. Here we observed both formal proofs of concrete security guarantees and empirical evaluations. While 3 security DSLs do not come with any of the two, one of them is an open standard and widely applied in practice. Therefore, we also investigated whether the security DSLs have been applied in practice, e.g., as part of a case study.
Guarantees and formal proofs
Overall, 40 security DSLs (33 %) state some form of security guarantee. However, there is a strong correlation between the availability of an attacker model and a security guarantee. We observed that if an attacker model is available, 79 % of these DSLs have a security guarantee and explicitly state the security objective. On the other hand, DSLs with a focus on Access Control do not have an attacker model (81 %) and do not provide a security guarantee (93 %) in most cases. If a security guarantee is provided by a security DSL, it should be proven that this guarantee holds under any conditions. Therefore, we investigated whether security DSLs are accompanied by formal proofs.
Not applicable.:
The majority of the security DSLs do not come with a formal proof since these are not applicable to them, e.g., due to their informal nature. Only 33 % of the DSLs provide a security guarantee, for example, many languages in the field of information flow control assert a form of non-interference, a concept introduced by Goguen and Meseguer (Goguen and Meseguer, 1982), which implies that the observable output of a system must be independent of its sensitive input. For the other 67 %, formal proofs are not applicable due to the lack of a security guarantee to provide a proof for.
Formal proof.:
Out of the set of DSLs with security guarantee, 73 % provide a formal proof that assures the correctness of the claimed guarantee. These are mostly from domains with strong formal foundations, e.g., cryptography or information flow control.
No proof.:
Unfortunately, 12 DSLs remain that contain security aspects that should be accompanied by a formal proof but for which no formal proof is provided.
Empirical evaluation
To get a better understanding how the security DSLs are evaluated, we first assessed using which empirical method the DSLs as well as supporting tooling are evaluated.
None.:
Unfortunately, most DSLs (44 DSLs) come without any empirical evaluation.
Example.:
A huge number of 29 DSLs illustrates the security DSL on smaller, typically to some extent realistic, yet still artificial examples. However, only for 17 DSLs these examples are actually part of an empirical evaluation. In these cases, the examples make up for a benchmark suite or are used in experiments with users of the DSL. In the other 12 DSLs, the examples are mainly a vehicle to illustratively explain the DSL to the reader but are sold as an empirical evaluation.
Experiment.:
Only for 16 DSLs their authors performed systematic experiments to evaluate the DSL, mainly concerning the aspect of performance (13 DSLs). Besides this, for 3 DSLs scalability is evaluated based on experiments and the effectiveness of 2 DSLs.
Case study.:
The main vehicle for empirical evaluation is the case study (42 DSLs). Thereby, the main objective of a case study is to assess the feasibility of the DSL for the intended use case in its natural context (Runeson and Höst, 2009). To this end, the DSL is applied to one or multiple case studies and discussed in the paper. On average, the papers in this category contain 2.5 case studies, with a maximum of 12 case studies, resulting in a median of 1 case study. The case studies themselves are real-world systems or realistic systems of varying sizes. External experts are involved only rarely.
User study.:
Seven user studies were conducted to evaluate the concrete usage of the security DSL by target stakeholders or in the concrete application context. The studies typically aimed to assess the usability for stakeholders (4 DSLs), the applicability to real-world contexts (4 DSLs), and the understandability (2 DSLs) and usability (1 DSL) of the DSL.
Benchmark.:
Four DSLs are evaluated by applying them to some sort of benchmark. Benchmarks are typically created for competitive evaluation in a specific domain and allow comparison of specific characteristics of the systems being evaluated (von Kistowski et al., 2015), in this case DSLs and their tools. However, only for two of them the entire benchmark suite is used, the other two selectively picked examples from benchmark suites.
Comparison.:
For one DSL a systematic comparison with other DSLs is provided in which the expressiveness of the DSLs is discussed based on their syntax.
111122223333444455556666777788889999101010101111111112121212131313131414141415151515161616161717171718181818191919192020202021212121222222220044448888121212121616161620202020Number of DSLsexamplescase-studies
Figure 11. Number of investigated case-studies/examples
Since examples and case studies are the main methods used for evaluating security DSLs, in Fig. 11, we depict the number of case-studies and examples presented for the DSLs. The maximum number is 22 for one DSL, but on average only 3.2 examples or case studies per DSL are presented. The median is 2 DSLs. Altogether, the empirical evaluation of the security DSLs is mostly focused on individual cases that may be considered representative, but rigorous evaluation methods are often missing.
After identifying the method used for evaluation, we investigated what are the objectives of the evaluations. In our investigation of the evaluation of the security DSLs we identified 13 different objectives of the evaluation as well as the illustration of the DSL on examples which is presented as an evaluation of the DSL.
Applicability.:
The applicability to real-world systems and application scenarios has been evaluated for 9 DSLs in the corresponding publications. Thereby, the focus is primarily on validating that variations that might occur in practice are supported technically and not whether the proposed technique provides good results for all of these variations.
Developer overhead.:
Only 1 DSL is evaluated concerning the overhead it introduces for developers using the DSL. However, this is assessed by the authors in a case study they performed and not by an user study as one would expect.
Effectiveness.:
The effectiveness to actually achieve the intended security goal is evaluated for 8 DSLs. Effectiveness is evaluated via controlled experiments. On average, 9.3 more or less representative examples are used for this purpose, with the minimum being only one example and the maximum being 22 examples.
Expressiveness.:
For 9 DSLs their expressiveness is evaluated. For this purpose, only what can be expressed using the syntax of the language is considered and not how the supporting tooling applies to real-world systems as also considered in the evaluations targeting applicability.
Feasibility.:
Feasibility is usually evaluated on individual case studies by demonstrating that the proposed technique is feasible to solving the targeted problem. Corresponding to case studies being the mainly applied evaluation method, the majority of the security DSLs (40 DSLs) is evaluated concerning their principle feasibility for the intended task.
Performance.:
The concrete performance of the security DSLs, typically focusing on the runtime of the corresponding tooling, is the evaluated for 18 DSLs.
Precision.:
For one DSL the precision of a static analysis based on information expressed using the security DSL is evaluated.
Scalability.:
For 4 security DSLs the evaluation systematically focuses in experiments on how the DSLs scale for huge inputs.
Understandability.:
Whether stakeholders can easily understand what is expressed using the security DSLs is evaluated for 2 DSLs.
Usability.:
For 5 DSLs the evaluation targets the concrete application of the DSLs by developers and how their experience in applying the DSL.
Usefulness.:
Whether the DSLs are useful in practice is evaluated for 2 DSLs. One time in a user study and once based on examples. Taking the benefits of the DSLs and supporting tooling as granted, the focus is on their embedding into practical processes and how the engineering of secure systems is improved in relation to overhead added.
Illustrate.:
The illustration of what the security DSL looks like to convey the idea of how it will be applied is based on examples for 12 security DSLs and is sold as a form of empirical evaluation. However, these examples are relatively small and may not fully align with the scope of a case study, which could potentially impact the validity of conclusions drawn from them.
Practical application
Only a minority of 14 security DSLs (12 %) has been applied in practice. Three of these DSLs are the OASIS standards SAML, XACML, and WS-SecurityPolicy that are widely used in practice. The mature academic security frameworks CORAS, STS, and UMLsec have been applied in various industrial context in projects and papers extending their original publications. All others have been applied to individual open-source projects or been evaluated together with industry partners.
Observation 6. Not only have most security DSLs rarely been applied in practice, their empirical validation is often lacking or of low quality, and even formal proofs are the exception. However, those DSLs that have been evaluated in practice or come from open source projects show a high interest from the industry.
5. Discussion
During our survey, we identified challenges related to the development and use of security DSLs as well as opportunities for future high-impact research. Overall, we developed a vision of the integrated utilization of security DSLs throughout the software development process as an effective means of developing secure systems. We first discuss identified challenges and opportunities, before outlining our vision and discussing why this integration will be an effective means.
5.1. Challenges & Opportunities
Although the use of security DSLs contributes to the security of a system, either by embedding security early on, i.e., during the requirements engineering or design phase, or by providing safe-to-use security measures during the implementation phase, the practical application seems to be limited. We identified few security DSLs that have been evaluated in practice, and even fewer DSLs that have become open standards. Investigating the barriers to practical adoption of security DSLs and the aspects that lead to successful adoption represents a significant research opportunity. The review of security DSLs and the findings from this survey are a first step in this direction.
Based on our extensive literature search, we can conclude that DSLs have already been presented for a wide range of security aspects and system components. We cannot identify a major security area where DSLs do not yet exist, but would obviously be helpful. While security DSLs exist for all major areas, there is still room for improvement in the details. For the access control domain, we observe that this detailed optimization is already well advanced, as DSLs have been explored for a variety of access control models. Also, for DSLs featuring information flow control, we see a trend toward more specific applications, such as concurrency or distributed systems.
Most DSLs have been published in security community venues, but note that the security community itself is fragmented into several subdomains. The isolation of each DSL ultimately reflects this fragmentation. Unfortunately, we observe a low level of integration and combination. Most of the languages we examined produce only single artifacts and are not integrated into other frameworks. Nevertheless, there are some security DSLs that cover multiple phases of the SDLC and demonstrate the benefits of doing so, i.e., ensuring that planned security requirements are addressed as planned in later development phases. Further integration promises a massive increase in the security of systems under development.
The lack of editing support we observed is probably typical of scientific implementations, which focus on novelty rather than usability. Nevertheless, a few security DSLs have led to commercial products, enjoyed widespread adoption beyond the research community, or become an international standard. However, effective editing support is needed to put the DSLs into practice, and involves relevant research opportunities, particularly for the human-based security community, such as identifying what developers need to effectively use a security DSL.
Particularly evident is, for scientific standards, the low level of evaluation for the DSLs we discovered. While the languages in the cryptography and information flow control domains usually provide strong formal security guarantees to undermine their effectiveness, most DSLs have not been evaluated with clear methodology. Many papers only demonstrate the feasibility of their language with small examples, while user studies have only been executed for a few DSLs. This is a call to arms to improve the evaluation of future work on security DSLs. However, this also provides an opportunity for empirical studies of the effectiveness of different security DSLs, including the potential to identify factors that influence the applicability or security gain of security DSLs.
5.2. Vision of an Integration Scenario
During our study, we identified various best practices and ideas that can help to effectively utilize security DSLs throughout the software development process. Based on these individual observations as well as the empirical data gathered, we developed a vision of deep integration of different security DSLs. To outline this vision and demonstrate how several of the DSLs examined in this work can be integrated to develop a secure system, we present a simple example of an EHR subsystem as motivated in the introduction. We map the DSLs to their respective phases in the SDLC, highlight how the DSLs help address specific security issues, and outline potential integrations. Our example consists of a measurement device, a computed tomography (CT) scanner, that sends its data over an open network to a server.
At the beginning of the development process, the (security) requirements are elicited and specified with the help of a security DSL, i.e., using STS and its supporting tools. The functional requirements of the system described above are modeled using a goal model notation oriented on SI* as shown in Fig. 6. Thereby, data planned to be exchanged between the individual entities is already explicitly specified on an abstract level. Here, the sensitivity of the measurement data gets identified and documented, for example, the CT images are labeled as security-critical data. Also using STS, basic authorization requirements for the intended usage scenarios are planned (see Fig. 12). For example, a doctor that is contacted as part of a remote consultation is allowed to read and modify health reports, while researchers are not allowed to have direct access. Thereby, checks provided as part of the editing support help in reasoning about the security requirements. For example, STS allows to propagate threats through the planned system to help identifying possible impacts. In Fig. 6, the tool shows by a red highlighting that the threat of the CT scanner being not reliable can propagate up to the treatment of a patient and might impact it.
[…]
After storing the data on the server, it should not be accessible to everyone and the authentication requirements shown in Fig. 12 must be realized. To this end, proper access control constraints must be specified so that, e.g., a physician can read the sensitive measurement values, but not the system administrator. Comparable to UMLsec, the early authorization requirements specified in STS should be taken as an input and be detailed using the more fine-grained access control DSLs. While we are not aware of any automation of the initialization of an access control DSL, DSLs such as UMLsec or SecureUML allow to immediately annotate design models with access control constraints, therefore supporting integrated development of more fine grained access control systems. In the best case, specification are created that can be transformed into a format that can be interpreted at runtime by an access control engine to avoid manual implementation of the access control policies.
To prevent unintended leakage of information also at a lower level, Jif is used to implement parts of the data gathering logic. For example, a measurement data variable can be declared in Jif as shown in LABEL:lst:jif. The security label in the curly brackets would assure, that the measurement data that is generated by the scanner can only be read by the scanner itself and any entity that can act as AuthorizedUser. Furthermore, the measurement can only be modified by the scanner to ensure its integrity. This code written in Jif can be compiled to Java and thus seamlessly integrated with code written in Java. However, as for the design models, the data flow policies expressed in Jif must comply to the data flows planned during requirements engineering (cf. Figs. 6 and 12) as well as all further refinements, e.g, using security DSLs at design time. While we are not aware of any compliance check between data flow policies in Jif and ones expressed in other security DSLs, there are works that check whether design-time data flow contracts are properly implemented using static or dynamic analysis or by tailoring taint analyzers (Tuma et al., 2023; Peldszus et al., 2024; Lanzinger et al., 2024)
Listing 4: Excerpt from a data class with secure data flow constraints written in Jif.
1public class CTScan {
2 // raw scan with data flow policy
3 byte[]{Scanner->AuthorizedUser; Scanner<-*} measurement;
4}
Finally, let us assume, the hospital participates in a medical study across multiple institutes, but privacy regulations prevent the disclosure of their patient’s measurement data. In response, secure multi party computation is applied, which allows computing functions, e.g., to determine average, median, and maximum values of the private data without revealing the data to other participating institutes. The involved parties can implement the functionality in SFDL 2.0 without cryptographic knowledge and the FairplayMP framework generates the desired code. As Jif, SFDL 2.0 compiles to Java bytecode that can be executed on the JVM, therefore allowing implementation-level integration.
Figure 12. Authorization view for an electronic health care system in STS.
The existence of several information-flow-sensitive DSLs, such as Jif or SFDL 2.0, points to the final challenge in the outlined integration, which is to decide on the appropriate DSLs to select for fulfilling the individual security requirements. Here, the overview provided in this paper helps to manually select suitable security DSLs for individual tasks. However, suitable security DSLs for the implementation of measures such as MPC should be automatically proposed as part of the development process, just as UMLsec was used to identify the required cryptography (cf. Fig. 8). Integrating such techniques into the reasoning of design-time security DSLs is the first step. The second step should be to automatically propose the most appropriate DSL that implements all the required features and best fits the design choices of the system.
6. Threats to Validity
To put our study into place, we discuss its validity concerning the four common threat categories of construct, external, internal, and conclusion validity (Jedlitschka et al., 2008). Overall, we are confident, that our methodology allowed us to identify all security domains, where a significant amount of DSLs have been published and that we include its most influential DSLs.
Construct Validity
A potential threat to the validity of this survey is that our methodology may have missed relevant domains that do not match our search terms and are not linked to, and thus referenced by, the papers we included. However, all DSLs known to us from our research experience were covered by this systematic approach, indicating that our search terms were chosen well.
The primary focus of the survey was put onto academic papers, although organizations in practice may be the main drivers of security DSLs. Due to this decision on the survey design, some media that are often used in practice, such as blogs discussing security DSLs, and that might better reflect practice, are not covered. However, since such works are not peer-reviewed, their quality is questionable and therefore excluded with good reason. Also, such practical applications are the subject of research that is explicitly covered by the search criteria used in this work.
Using Semantic Scholar as the literature database for the search may affect the works that can be found. However, Semantic Scholar is a database that aggregates results from all major publishers. Therefore, the impact of possible threats due to the choice of Semantic Scholar is minimal.
Internal Validity
There might be a bias due to the research background of the authors conducting the study, which could have impacted the execution of the study, e.g., decisions on what is considered as a security DSL. To mitigate this threat, two authors with different backgrounds, the first author with a background in cryptography and side-channel attacks and the second author in security by design, conducted each step of the investigation.
A second internal threat could arise from possible misunderstandings of the reviewed papers by the authors, which could lead to incorrect classifications. To mitigate this threat, in cases where the authors were unsure about a classification, the authors actively discussed these classifications, with other authors reading the papers in question. In addition, when analyzing the results, the authors actively read papers where the extracted statements and corresponding classifications did not seem to match. However, only minor discrepancies were identified, all of which were borderline, indicating a good quality of the extracted classifications.
External Validity
The main threat to external validity is that the papers analyzed may not be representative of security DSLs applied in practice or of all security DSLs considered in the security research community. However, the significant number of papers examined and security DSLs extracted reduces the potential impact of this threat. Furthermore, the security DSLs examined come from many different research communities and cover both scientific and practitioner-driven security DSLs, which indicates a good generality of the results.
Conclusion Validity
In practice, DSLs may be more deeply integrated than observed but not yet investigated by the research community. Therefore, this could be considered the main threat to the conclusions drawn from this study. However, since the statistics of the examined security DSLs show that most of them are relatively old, but our search also covers newer works, this is unlikely.
7. Related Work
Only a few works have systematically examined security-related DSLs, and most of them have a narrower scope than this survey. Two existing works (do Nascimento et al., 2012; Kosar et al., 2016) give an overview of the domains, where DSLs have been used, but are not security specific. On the other hand, multiple surveys review tools, frameworks and DSLs for single security subdomains. Erata et al. (Erata et al., 2023) review tools for formal verification of security properties in software and hardware and include domain-specific hardware description languages for secure information flow. Tools and DSLs for the development, specification and implementation of cryptography have been evaluated in the context of computer-aided cryptography (Barbosa et al., 2021). Hastings et al. (Hastings et al., 2019) systematize the very specific cryptographic subdomain of multi party computation. Another area, that induced many new DSLs in the last years are smart contract languages and this area has been contextualized by two recent works (Varela-Vaca and Quintero, 2022; Vacca et al., 2021). A subset of the DSLs that we cover are included in publications specifically about policy specification languages (Han and Lei, 2012; Kasem-Madani and Meier, 2015; Leicht and Heisel, 2019). Also, from the domain of security-aware system modeling, DSLs have been surveyed (Hachem et al., 2016). Lallie et al. (Lallie et al., 2020) review the visual syntax of attack graphs. We identified four surveys (Uzunov et al., 2012; Nguyen et al., 2015; van Den Berghe et al., 2017; Mashkoor et al., 2023) on model-driven security engineering that mention various design-time security DSLs but mainly focus on model-based security engineering processes. A systematic overview and comparison, such as in this work, is not provided.
8. Conclusion
With this work, we present the first holistic survey of domain-specific languages for security, by analyzing 120 security DSLs based on six research questions. In particular the security research community already proposed security DSL for a wide range of security aspects. Most DSLs, address the design or implementation phases by enhancing the expression of design models, policies or code. We conclude that the integration, combination, usability, and evaluation of security-oriented DSLs still offers many opportunities for further research.